xx SystemTopology

SystemTopology

Erik Schnetter <schnetter@gmail.com>

April 26 2015

Abstract

1 Introduction

2 Physical System 3 Numerical Implementation 4 Using This Thorn 4.1 Obtaining This Thorn 4.2 Basic Usage 4.3 Special Behaviour 4.4 Interaction With Other Thorns 4.5 Examples 4.6 Support and Feedback 5 History 5.1 Thorn Source Code 5.2 Thorn Documentation 5.3 Acknowledgements References 6 Parameters

11 setxthreadxbindings10 Scope: private01  KEYWORD

11 Description: Set thread bindings (aka thread-CPU affinity)

11 Range 10 01 Default: auto
11 yes  10 set bindings
11 no  10 no not modify bindings
11 auto  10 depends on system architecture

11

11 threadxlayout  10 Scope: private01  KEYWORD

11 Description: How to lay out threads over cores

11 Range 10 01 Default: dense
11 dense  10 Pack threads on as few NUMA nodes/cores as possible
11 loose  10 Spread threads over all NUMA nodes/cores

11

7 Interfaces

General Implements: systemtopology

Provides: GetNumSMTThreads to GetMaxSMTThreads to GetCacheInfo1 to GetMPIProcessInfo to

8 Schedule This section lists all the variables which are assigned storage by thorn CactusUtils/SystemTopology. Storage can either last for the duration of the run (Always means that if this thorn is activated storage will be assigned, Conditional means that if this thorn is activated storage will be assigned for the duration of the run if some condition is met), or can be turned on for the duration of a schedule function.

Storage NONE

Scheduled Functions

CCTKxSTARTUP x stxsystemxtopology x output and/or modify system topology and hardware locality x

11 x 11 After:  11 hwlocxstartup
11 x 11 Before: 11 driverxstartup
11 x 11 Language: 11 c
11 x 11 Type:  11 function